On March 10, 2019, Lost World Company sells equipment that it purchased for $192,000 on August 20, 2012. It was originally estimated that the equipment would have a life of 12 years and a salvage value of $16,800 at the end of that time, and depreciation has been computed on that basis. The company uses the straightline method of depreciation.

Following are the assumptions with respect to partial periods:

(1) Depreciation is computed for the exact period of time during which the asset is owned. (Use 365 days for the base and record depreciation through March 9, 2019.)
(2) Depreciation is computed for the full year on the January 1 balance in the asset account.
(3) Depreciation is computed for the full year on the December 31 balance in the asset account.
(4) Depreciation for one-half year is charged on plant assets acquired or disposed of during the year.
(5) Depreciation is computed on additions from the beginning of the month following acquisition and on disposals to the beginning of the month following disposal.
(6) Depreciation is computed for a full period on all assets in use for over one-half year, and no depreciation is charged on assets in use for less than one-half year. (Use 365 days for base.)


Briefly evaluate the methods above, considering them from the point of view of basic accounting theory as well as simplicity of application.
Answer #1

Cost = $192000
Life = 12 years
Salvage value = $16800
Depreciation under straight line = [(Cost - Salvage value)/ Life]
=> [($192000 - $16800)/ 12 years] = $14600 per year

(1)
In 2012, Machine purchased on August 20. So number of days = 134 days till 31 december 2012
therefore depreciation = $14600 X (134 days / 365 days) = $5360
Depreciation for 2013-2018 for entire year = $14600 X 6 years = $87600
In 2019, machine sold at March 10, So number of days from January 01 = 68 days
therefore depreciation = $14600 X (68 days / 365 days) = $2720
So Total depreciation = ($5360 + $87600 + $2720) = $95680

(2)
In the asset account for January 1 balance it will be computed for the whole year. In 2012 it was purchased in month of August meaning that January 01, 2012 it did not exist in Asset account. So no depreciation for 2012.
In 2019 it was sold in March month showing that it did exist in January 01, 2019 asset's account. So full yeaar depreciation for 2019.
Thus depreciation will be charged from 2013-2019 = 7 years
So, Depreciation = $14600 X 7 years = $102200

(3)
Using the same concept as applied above in (2). It was purchased before December month in 2012 so it did exist in December 31,2012 asset's account but was sold before December 31, 2019.
Therefore Depreciation will be charged from 2012-2018 = 7 years
So, Depreciation = $14600 X 7 years = $102200

(4)
So basically for year of acquisition and sale, half year depreciation will be charged and for other full year depreciation
=> Depreciation = (2012 - Half) + (2013 to 2018 - Full) + (2019 - Half)
=> ($14600 X 1/2 year ) + ($14600 X 6 years) + ($14600 X 1/2 year)
=. $102200

(5)
Meaning for the month of acquisition we will assume depreciation from September 1 instead of purchase date of August 20 = 4 months till December 2012
So, Depreciation for 2012 = ($14600 X 4/12 months) = $4866.67
For 2013-2018 full years = ($14600 X 6 years) = $87600
For month of sale, we will assume depreciation till April 01, 2019 instead of actual selling date of March 10, 2019 = 3 months since January 2019
So, Depreciation for 2019 = ($14600 X 3/12 months) = $3650
Total depreciation = ($4866.67 + $87600 + $3650) = $96116.67

(6)
Half of total year = (365 days/2) = 182.5 days
In (1) above we calculated that number of days for the year of acquisition = 134 days and number of days for year of sale = 68 days. Since both are less than 182.5 days, therefore no depreciation for year of acquisition and sale
Depreciation for 2013-2018 = ($14600 X 6 years) = $87600